Calculate the profit under 1st-degree price discrimination for a monopolist facing the demand curve P = 12 − Q and cost given by T C = 5 + 2Q + 1.5Q^2

.
Answer #1

Answer

The firm produces at MC=P

MC=change in total cost =first differentiation of the TC function =2+3Q

equating MC=P

2+3Q=12-Q

4Q=10

Q=2.5

P=12-2.5=9.5

=====

The profit is equal to the consumer surplus for the firm as the firm charges price equal to the willingness to pay of each consumer and the willingness to pay is depicted by the demand curve, the CS is below demand and above price

CS=profit =0.5*( Y-axis intercept fo the demand curve -P)*Q

=0.5*(12-9.5)*2.5

=$3.12500

The profit is $3.125